### v1

# Learning Selection Cuts With Gradients

Abstract
Many analyses in high-energy physics rely on selection thresholds (cuts) applied to detector, particle, or event properties. Initial cut values can often be guessed from physical intuition, but cut optimization, especially for multiple features, is commonly performed by hand, or skipped entirely in favor of multivariate algorithms like BDTs or classification networks. We revisit this problem, and develop a cut optimization approach based on gradient descent. Cut thresholds are learned as parameters of a network with a simple architecture, and can be tuned to achieve a target signal efficiency through the use of custom loss functions. Contractive terms in the loss can be used to ensure a smooth evolution of cuts as functions of efficiency, particle kinematics, or event features. The method is used to classify events in a search for Supersymmetry, and the performance is compared with traditional classification networks. An implementation of this approach is available in a public code repository and python package.
1 Introduction
Rule-based data analysis is a common starting point for many studies in high-energy physics. Specific physics processes often motivate phase space requirements, such as cuts on particle momenta or event activity, that are based on physical principles. Optimization of those cut values, using significance or other metrics, can lead to event selections that are easily understood in the context of the physics signature under study. While an analysis based only on one-dimensional cuts will rarely match the performance of a multivariate optimization that fully exploits correlations between particle or event features, the transparency of cut-based procedures, and the relatively straightforward interpretations of the phase space captured by the cuts, make them attractive as a starting point for data analysis.
Cut-based event selections are also useful in other ways. A set of one-dimensional cuts constructs a contiguous volume in feature space that is naturally robust against training samples that do not fully span all features, while deep neural networks often identify islands in feature space whose definitions can be more sensitive to the details of the training procedure. This can be especially undesirable in cases where the training data are derived from samples of simulated events that may not perfectly model the event features used during evaluation. The signal efficiency of a single volume in feature space may be easier to calibrate in such cases, which can be useful in measurements where excellent control over signal efficiency is required. Cut-based selections are also commonly used in event generators and hardware triggers, where more complicated selections based on neural networks are challenging to implement (though these challenges are quickly being overcome, see [Jia:2024ysq] for a recent review of machine learning in hardware).
Despite the utility of cut-based approaches, tools to derive optimal cut values are limited, and mostly consist either of TMVA’s kCuts routine [hoecker2009tmvatoolkitmultivariate], or brute-force scans of feature space (see e.g. the optimization package). The kCuts algorithm is a common starting point for cut-based analyses, and derives cuts that maximize the background rejection at different signal efficiency working points. Our work was partially motivated by the observation that kCuts tunes the cuts for each target efficiency separately, such that small changes in the target efficiency can result in large changes in cuts. Subsets of events with similar (but not identical) features can also lead to very different cut values when treated separately in training. We desired an algorithm that was relatively insensitive to the exact value of the target signal efficiency, and that could provide well-behaved (ideally monotonically-varying) cuts as functions of common event properties like pileup or particle momenta. This also presented an opportunity to implement such an algorithm using modern machine learning libraries.
This work presents an approach to binary classification where optimized cuts are extracted from the parameters of a trained neural network with an extremely simple architecture. Training is performed using gradient descent, requiring the modeling of cuts using differentiable functions. The application of a set of selection criteria is modeled as a product of sigmoid-activated linear transformations of input features, where the biases and weights of the linear transforms can be translated into cuts. We refer to this as the "cuts as biases in networks" (CABIN) approach. Custom loss functions are used to derive cuts for specific target efficiencies, and to regulate the evolution of cut values across efficiency targets and other event properties. We test the performance of this technique using a randomly generated dataset, and study the stability of the training procedure for different network hyperparameters. We also investigate the performance of this method in a study of LHC sensitivity to SUSY models using samples of simulated events. Ref. [Watts:2024wes] studies this problem in the context of cuts used as preselection requirements prior to the use of a neural network classifier, and confronts some of the challenges we discuss here, such as the expression of cuts as differentiable functions. Our work focuses on the optimization of cuts, without including any additional multivariate classifiers, and introduces contractive terms to control the behavior of cuts across working points.
2 Learning cuts as biases
We consider the problem of assigning binary labels to events $\mathbf{x_{i}}$ of a set of events. Each $\mathbf{x_{i}}$ can be represented as a vector of input features: $\mathbf{x_{i}}=(x_{i,1},...,x_{i,k},...,x_{i,m})$ where we use $k$ to index the features from $1$ to $m$, and $i$ to index the events.
In a typical neural network trained for binary classification, the $m$ features are treated as inputs for a fully-connected feed-forward network that has a single output, representing a sum of linear transformations of the input values passed through one or more non-linear activation functions, that provides a score for a given event $\mathbf{x_{i}}$. After training, events from one category will cluster at one end of the network output score distribution, while events from the other category cluster at the other end of the range of scores. A threshold on the output score distribution can then be chosen based on signal efficiency, background rejection, significance, or other metrics. In this work we are interested in deriving thresholds on the input features, not on the summary score.
We apply a linear transformation to each input feature $x_{i,k}$:
$x_{i,k}^{\prime}=w_{k}x_{i,k}+b_{k}$ (1)
where $w_{k}$ represents a weight, and $b_{k}$ is the bias. Here we consider a simple case where positive $x_{i,k}^{\prime}$ correspond to one category, and negative values correspond to another category, with threshold value that separates between the two at $x_{i,k}^{\prime}=0$. The threshold in the untransformed space is then:
$x_{i,k}<-\frac{b_{k}}{w_{k}}.$ (2)
If the weights are chosen to be unity (up to a sign) then the bias $b_{k}$ represents the cut value for feature $k$. As a parameter of a linear layer in a model, the optimal value of the cut can be learned through training.
2.1 Differentiable cuts
The problem with learning a set of rectangular cuts, which is equivalent to a product of Heaviside step functions $\Theta_{k}(x)$, is that step functions cannot be used for gradient-based training. We address this by approximating $\Theta_{k}(x)$ with the logistic function, $\sigma(x)$, where the inputs are multiplied by some factor $n$:
$\Theta(x)=\lim_{n\to\infty}\sigma(nx)=\lim_{n\to\infty}\frac{1}{1-e^{-nx}}$ (3)
For finite $n>1$, this function is differentiable and provides a good approximation of the Heaviside function as long as $x$ is not near 00, the threshold used to distinguish between categories. Figure 1 shows the behavior of the logistic function for different values of $n$.
Figure 1: The logistic function $\sigma(x)$ can be used to approximate the Heaviside step function $\Theta(x)$ by scaling the inputs. This provides a differentiable proxy for rectangular cuts.
The result of applying all cuts on features $1..m$ for event $i$ is then calculated as the product of the individual cut functions:
$\mathrm{score}_{i}=\prod_{k=0}^{m-1}\sigma(nx_{i,k}^{\prime})$ (4)
The score goes to zero if the event fails any cut, and only converges to a value near one if the event passes all cuts, and by a sufficient margin that they are away from the sigmoid turn-on for every input feature.
The final output of this algorithm is a score, not a binary decision. The scores asymptotically approach 00 or $1$ for events that fail or pass the selection cuts, respectively, and for sufficiently large values of $n$ the sum of scores for events that are evaluated by the network closely matches the number of events that pass the cuts learned in the training procedure. On the other hand, values of $n$ that are too large can make learning unstable, as gradients vanish quickly when $x^{\prime}$ values move away from zero.
2.2 Loss Functions
The usual loss function used for two-category classification is binary cross entropy (BCE), which penalizes incorrect classification of any objects in the dataset. Another approach, used by the kCuts algorithm, is to tune the cuts to yield a specified target efficiency, and then maximize the background rejection for that efficiency. In order to reproduce the TMVA procedure, which is a common starting point for many HEP analyses, we implement a loss function, referred to here as the "efficiency loss", based on deviations from a target signal efficiency, in addition to penalizing non-zero background efficiency:
$\mathrm{loss}_{\mathrm{efficiency}}=\alpha\left(\varepsilon_{\mathrm{signal}}^{\mathrm{target}}-\varepsilon_{\mathrm{signal}}^{\mathrm{estimated}}\right)^{2}+\beta\varepsilon_{\mathrm{background}}^{\mathrm{estimated}}$ (5)
where $\varepsilon_{\mathrm{signal}}^{\mathrm{estimated}}$ is the sum of output scores for true signal events, normalized to the number of true signal events, and $\varepsilon_{\mathrm{background}}^{\mathrm{estimated}}$ is the sum of output scores for true background events, normalized the number of true background events. Positive-valued hyperparameters $\alpha$ and $\beta$ can be used to control the relative importance of signal efficiency vs background rejection during training. As discussed in Section 2.2, the choice of $n$ in equation 4 influences how closely $\varepsilon_{\mathrm{signal}}^{\mathrm{estimated}}$ approximates the actual signal efficiency obtained by applying the cuts derived from the network parameters. Since $\varepsilon_{\mathrm{background}}^{\mathrm{estimated}}\geq 0$, the loss defined in equation 5 is convex.
Implementation of a custom efficiency loss function also allows the introduction of contractive terms that regulate the behavior of cuts individually and across different optimization regions. The first contractive term we consider is based on the magnitude of the cuts themselves:
$\mathrm{loss}_{\mathrm{cut\ size}}=\frac{\gamma}{m}\sum_{k=0}^{m-1}b_{k}^{2}$ (6)
where $b_{k}$ are the biases, $m$ is the number of features, and $\gamma$ is a positive-valued hyperparameter that controls the influence of this loss term during training. This contractive term provides a penalty based on the magnitude of the cut that avoids having cut values that are far from the training distributions.
To enable easy comparisons of the BCE loss with custom loss terms, a BCE loss term, scaled with parameter $\delta$, is also added to the efficiency loss. The $\delta$ parameter can be set to small values (or zero) to ignore the BCE term, or can be set to one with all other loss hyperparameters set to zero to produce exactly the BCE loss.
3 Efficiency scan networks
The one-to-one linear networks described above, in combination with the custom loss function that trains a network to yield a target signal efficiency, allows the construction of a collection of simultaneously-trained networks that span multiple efficiency targets. An associated loss function, called the efficiency loss, incorporates efficiency loss terms for each of the sub-networks, as well as contractive terms that regulate the behavior of the cut values across different sub-networks. The contractive terms are implemented as penalties for cuts that are not between cuts for neighboring target efficiencies:
$\mathrm{loss}_{\mathrm{smooth}}=\varepsilon\sum_{k=1}^{m-2}\frac{\left(b-\bar{b}\right)^{2}}{(\Delta b)^{2}+\rho}$ (7)
where $b$ is the bias (cut) for a given target efficiency, $\bar{b}=\frac{b_{+}+b_{-}}{2}$ is the mean of the biases for adjacent higher ($b^{+}$) and lower ($b^{-}$) target efficiencies, respectively, and $\Delta b=b^{+}-b^{-}$ is the range between adjacent cut values. This loss term encourages the cut to be between adjacent cuts, while $\rho$, which we set to 0.1, avoids divergences in cases where the cuts for adjacent target efficiencies are identical (e.g. at the beginning of training).
4 Implementation
We implement the CABIN method in the cabin python package [cabin] using pytorch [pytorch]. A simple, one-to-one linear network is constructed in which $m$ input features feed to $m$ nodes in a single hidden layer, with each input connected to one and only one hidden node. The network is illustrated in Fig. 2. The input feature first undergoes the linear transformation in equation 1, and then goes through a logistic activation function. The outputs of the hidden layer are multiplied to produce a single output score for each event:
$\mathrm{score}_{i}=\prod_{k=0}^{m-1}\sigma\left(n\left(w_{k}x_{i,k}+b_{k}\right)\right)$ (8)
We find that $n=15$ results in output scores for signal events that are close enough to unity that the sum of the output scores is a good estimate of the number of signal events that pass cuts derived from the weights and biases of the network. The weights can either be fixed by the user, or learned as part of the training procedure. We prefer setups in which the weights $w_{k}$ are fixed to ±$\pm 1$, such that they only encode whether the cut represents a "greater than" or "less than" requirement.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$h_{1}=\sigma(w_{1}x_{1}+b_{1})$$h_{2}=\sigma(w_{2}x_{2}+b_{2})$$h_{3}=\sigma(w_{3}x_{3}+b_{3})$$h_{4}=\sigma(w_{4}x_{4}+b_{4})$$y=\prod h_{k}$$b_{1}$$b_{2}$$b_{3}$$b_{4}$
Figure 2: A one-to-one linear network used to learn cuts as biases. The version illustrated here has fixed weights, which determine whether the associated bias corresponds to a "greater-than" or "less-than" cut.
The training procedure is sensitive to the initial values assigned to the weights and biases. If the weights and biases result in cuts that are far from the bulk of the distribution then the gradients of the loss function can vanish, leading to poor training convergence. Initializing cut values in the middle of the distribution helps with this. We use the StandardScaler function in scikit-learn [scikit-learn] to rescale the inputs to have zero mean and unit width. The biases are then initialized to 0, which helps avoid vanishing gradients at the start of training.
For comparison with more traditional approaches, we also implement a simple version of a linear network, with $m$ inputs, one output, and no hidden layers, illustrated in Figure 3. The network output is calculated as:
$\mathrm{score}_{i}=\sigma\left(\sum_{k=0}^{m-1}w_{k}x_{i,k}\right).$ (9)
This network, with trainable weights and no biases, has the same number of free parameters ($m$) as the one-to-one linear network in equation 8 that has trainable biases and fixed weights. A BCE loss function is used to evaluate this model.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$y=\sigma\left(\sum_{k}w_{k}x_{k}\right)$$w_{1}$$w_{2}$$w_{3}$$w_{4}$
Figure 3: A simple linear network, with trainable weights and no bias term, used for comparison with the one-to-one linear network that learns cuts as biases.
Networks are trained with the torch.optim.SGD optimizer with a learning rate of 0.5, processing data in a single batch for 100 to 200 epochs. Since these networks are simple by construction, training times are typically fast, and no attempts were made to optimize the training hyperparameters or tune for fast convergence of the loss. The hyperparameters included in the loss functions are summarized in Table 1, along with values used during training.
Parameter Use Typical value
$\alpha$ Importance of reaching target signal efficiency 1e+1
$\beta$ Importance of minimizing background 1e-1
$\gamma$ Overall magnitude of cuts 1e-5
$\delta$ BCE term 1e-3
$\varepsilon$ Smoothness of cuts across signal efficiencies 1e0
Table 1: Hyperparameters used in the custom loss functions. In studies of the BCE loss, $\delta$ is set to 1.0 and all other terms are set to 0. When deriving cuts for target efficiencies, $\delta$ can be zero or small compared to $\alpha$, $\beta$, and $\epsilon$.
5 Data
We use a dataset of 20,000 events with ten features, where event properties are generated by randomly selecting values from Gaussian distributions with randomly-chosen means and widths. No correlations between features are implemented. Feature distributions for this dataset, referred to here as the Gaussian dataset, are shown in Fig. 4. We use 80% of the Gaussian dataset for training, and reserve 20% for testing.
Figure 4: Feature distributions in the Gaussian dataset for signal (red) and background (blue) events used for initial development of the CABIN approach. Features are Gaussian-distributed with assorted means and widths. No correlations between features are implemented.
A second dataset is used to study the behavior of the CABIN approach in a collider physics context. We consider the problem of discriminating events arising from Supersymmetric interactions from those of the Standard Model (SM). In particular, we simulate the production of pairs of sleptons, scalar SUSY partners of SM leptons, which are separated in mass from the lightest neutralino, here also the lightest SUSY partner, by approximately 20 GeV. The final state includes two opposite-sign leptons from the slepton decays, along with missing transverse energy ($E_{\mathrm{T}}^{\mathrm{miss}}$) from the invisible neutralinos. In such scenarios, it is often advantageous to require a hard jet, nominally from initial state radiation, which boosts the SUSY system. The SM production of $WW$ events, with both bosons decaying leptonically, provides an irreducible background that is difficult to suppress. Example Feynman diagrams of the SUSY signal and irreducible SM background process is shown in Fig. 5.
Figure 5: Feynman diagrams for SUSY slepton production (left) and background from SM diboson production (right).
We generate events for signal and background processes using MadGraph5_aMCNLO version 5.2.7.3 [Alwall:2014hca] using the NNPDF 2.3 leading order parton distribution functions [Ball:2012cx]. Events are showered with Pythia version 8.244 [Sjostrand:2014zea], and passed through a parameterized simulation of the ATLAS detector response implemented in Delphes version 3.5.0 [deFavereau:2013fsa, Selvaggi:2014mya, Mertens:2015kba]. Events in the SUSY search are pre-selected to have two leptons, at least one energetic jet ($p_{\mathrm{T}}>100$ GeV), and substantial missing transverse energy ($E_{\mathrm{T}}^{\mathrm{miss}}>300$ GeV). After preselection, we consider the following event features:
  • $E_{\mathrm{T}}^{\mathrm{miss}}$
  • the $p_{\mathrm{T}}$ of the leading jet
  • the separation in $\phi$ between each lepton and the missing transverse momentum: $\Delta\phi(\ell,E_{\mathrm{T}}^{\mathrm{miss}})$
  • the transverse mass of each lepton and the $E_{\mathrm{T}}^{\mathrm{miss}}$:
    $m_{\mathrm{T}}=\sqrt{2E_{\mathrm{T}}^{\mathrm{miss}}p_{\mathrm{T}}(\ell)\left(1-\cos\Delta\phi\left(\ell,E_{\mathrm{T}}^{\mathrm{miss}}\right)\right)}$
  • the separation $\Delta R$ of the leptons from each other:
    $\Delta R=\sqrt{\left(\Delta\phi(\ell_{1},\ell_{2})\right)^{2}+\left(\Delta\eta(\ell_{1},\ell_{2})\right)^{2}}$
Figure 6: Distributions of kinematic features used to classify SUSY and SM $WW$ events.
6 Results
We first study the behavior of several network configurations using the Gaussian dataset. We consider: a traditional linear network with a BCE loss function; a one-to-one linear network with a BCE loss function; a one-to-one linear network with an efficiency loss function; and an ensemble of one-to-one linear networks trained at different efficiency working points. We then study the ability of these different networks to classify simulated LHC events as arising from Standard Model interactions or from production of new particles in the Minimal Supersymmetric Standard Model (MSSM).
6.1 Gaussian dataset
The simple linear network described in section 4 is trained using the Gaussian dataset, with results shown in Figure 7. Despite the simple network architecture, the performance of this network is very strong, with an AUC of over 0.99. A zero-background, high-efficiency region can be constructed through cuts of around 0.8 on the output score. While more complex models could likely improve on these results, this already represents near-perfect classification, and will serve as a basis for comparison with other network structures.
Figure 7: Results of training the fully-connected network with a BCE loss function and the Gaussian dataset. The output scores (left) show good separation of signal and background for the Gaussian-distributed features in the generated dataset, even with the relatively simple network architecture. The ROC curve and AUC (right) provide a measure of the performance of the network.
The one-to-one linear network is trained in the same way, using BCE loss, with results shown in Fig. 8. The classifier is not as strong, since this model is equivalent to applying cuts on all features such that signal events failing any one cut are misclassified, while the simple linear network can compensate for background-like behavior in one feature with signal-like behavior in other features. The biases of the one-to-one linear network can be translated into cuts on the input features, shown in Fig. 9. The signal efficiency (true positive rate) of these cuts applied to the test sample is 93%, while the background efficiency (false positive rate) is 5%.
Figure 8: The output scores (left), ROC curve and AUC (right) after training the one-to-one linear network with a BCE loss function and the Gaussian dataset.
Figure 9: Gaussian dataset features for signal (red) and background (blue) shown with cuts optimized using a BCE loss function. Dashed lines represent "greater-than" cuts and dotted lines represent "less than" cuts. The cuts have a signal efficiency of 93% and a background efficiency of 3%.
The one-to-one linear model is next trained with the efficiency loss rather than BCE loss, where we set the target efficiency to be 93% to match the output of the previous configuration. After training, the cuts extracted using the efficiency loss give an actual signal efficiency of 92% with a background efficiency of 2.7%, with nearly identical cuts to the network trained with BCE loss. The choice of hyperparameters $\alpha$ and $\beta$, which control the importance of reaching the target signal efficiency and minimal background efficiency, respectively, influences how close the network gets to the target signal efficiency before the background efficiency becomes too large. In this case, the target signal efficiency is not reached, in favor of keeping the background efficiency small. Some tuning of the $\alpha$ and $\beta$ parameters will likely be needed for other applications depending on the composition of the dataset and the desired signal efficiency or background rejection.
Finally, an ensemble of one-to-one linear networks trained with the contractive efficiency loss function to create cuts that target signal efficiencies from 20% to 90% in steps of 10%. Additional models for target signal efficiencies of 93% and 95% are also included, to investigate the regime where background rejection and signal efficiency begin to compete in the minimization of the loss.
The loss distribution for the efficiency scan network is shown in Fig. 10, including the contributions from the different terms in the training loss. The loss terms associated with reaching the desired signal efficiency and minimizing the background efficiency dominate the overall loss, but the smaller terms associated with the size of the cuts and the smoothness of cuts across target efficiencies regulate the behavior and prevent the cuts from settling into inappropriately loose or tight values. The impact of the BCE term is negligible, and removing it has no impact on the derived cuts in this fit configuration.
Figure 10: The loss distribution for the efficiency scan network, for training and test subsets of the Gaussian dataset. The loss is composed of different terms, corresponding to the signal efficiency, background efficiency, cut sizes, and smoothness conditions, along with a BCE term, all scaled by independent hyperparameters listed in Table 1.
The output score distributions and ROC curves for the trained networks are shown in Figs. 11 and 12. The ROC curves show that the networks trained for low signal efficiencies are far from optimal as measured by AUC. The score distributions show that the signal peaks near zero and near one, with events migrating from zero to one as the target efficiency rises. Due to the way the nets are trained, the vast majority of events passing the trained cuts have output scores above 0.5, and all events that fail cuts have output scores less than 0.5. Taken together, this implies that for a given target efficiency, the effective cut on the output score distribution that corresponds to an application of the trained cuts is near 0.5, in a part of the output score distribution that is not rapidly changing. This makes the calculation of the output score less sensitive to variations in the input features that might lead to changes in the output score. For comparison, the output score corresponding to 40% (50%) for the network with BCE loss is 0.9999994 (0.9999969), so changes of a few parts in $10^{-6}$ is enough to change the efficiency in such a network.
Figure 11: Output score distributions for different target efficiencies in the efficiency scan ensemble trained using the Gaussian dataset.
Figure 12: ROC curves for different target efficiencies in the efficiency scan ensemble trained using the Gaussian dataset.
Figure 13 shows the evolution of cut values across different target efficiencies, with a comparison to the BCE-trained cuts at an efficiency of 93%. The cuts vary smoothly across efficiencies, as desired, and closely match the cuts for the BCE-trained net. The tension between reaching the target signal efficiency and minimizing the background efficiency, especially at large target efficiencies, is shown in Fig. 14.
Figure 13: Cuts on Gaussian dataset features extracted from the trained one-to-one linear networks in the efficiency scan ensemble, as a function of the target efficiency. The cuts from the BCE-trained network are shown for comparison.
Figure 14: Signal and background efficiencies as a function of the target efficiency for the Gaussian dataset, both calculated by the efficiency loss function and after applying the cuts post-training, for subnets in ensemble of one-to-one linear networks.
6.2 SUSY dataset
The set of SUSY and SM $WW$ events is used to study the CABIN approach in a more realistic context for collider physics, in which the collection of features does not provide perfect discrimination between signal and background. The seven features described in Section 5 are used as inputs. Results of training the simple fully-connected network on the SUSY dataset using a BCE loss function are shown in Fig. 15. This configuration is the only case in which the learning rate was changed from 0.5 to 0.1, to avoid oscillations observed in the loss when near the minimum. The simple fully-connected network reaches an AUC of 0.69, reflecting the increased difficulty of this classification task relative to classification in the Gaussian dataset. [^1] For comparison, a network with two hidden layers and 20 nodes/layer achieves an AUC of 0.85, on the same dataset. This is further seen in Figs. 16 and 17, where the efficiency scan network is able to achieve higher ROC scores than the fully-connected network for all target efficiencies. In all cases, the score distribution is bimodal, reflecting the binary pass/fail decisions the network is modeling.
Figure 15: Results of training the fully-connected network with the SUSY dataset.
Figure 16: Output score distributions for different target efficiencies in the efficiency scan ensemble when trained with SUSY dataset.
Figure 17: ROC curves for different target efficiencies in the efficiency scan ensemble trained using the SUSY dataset.
The loss distribution for the efficiency scan network trained on the SUSY dataset is shown in Fig. 18, including the contributions from the different terms in the training loss. The loss terms associated with reaching the desired signal efficiency and minimizing the background efficiency dominate the overall loss, but the smaller terms associated with the size of the cuts and the smoothness of cuts across target efficiencies regulate the behavior and prevent the cuts from settling into inappropriately loose or tight values. The impact of the BCE term is negligible, and removing it has no impact on the derived cuts in this fit configuration.
Figure 18: The loss distribution for the efficiency scan network, for training and test subsets of the SUSY dataset. The loss is composed of different terms, corresponding to the signal efficiency, background efficiency, cut sizes, and smoothness conditions, along with a BCE term, all scaled by independent hyperparameters listed in Table 1.
The optimized cuts on the SUSY dataset for the efficiency scan network and for a one-to-one linear network trained with a BCE loss function are shown in Fig. 19. The network trained with a BCE loss function has a signal efficiency of 67% and a background efficiency of 41%, and fails to converge on cuts that perform as well as the efficiency scan network. This indicates that the smoothing terms in the loss function do not always lead to worse performance, and can help to avoid convergence to a local minimum that the learning routine cannot escape. The efficiency scan network also succeeds at ensuring a smooth (but still nonlinear) evolution of cut values for the individual input features in the SUSY dataset.
Figure 19: Cuts on SUSY dataset features extracted from the trained one-to-one linear networks in the efficiency scan ensemble, as a function of the target efficiency. The cuts from the BCE-trained network is shown for comparison.
The correspondence between the approximated efficiencies computed in the loss function and the actual efficiencies of the trained cuts is shown in Fig. 20. Here it is apparent that the background efficiency rises proportionally with the signal efficiency, rather than remaining zero for all but the highest target signal efficiencies as with the Gaussian dataset. The choice of a final working point for an analysis would depend on factors such as the statistical and systematic uncertainties on the background yields and the signal cross section.
Figure 20: Signal and background efficiencies, calculated in the efficiency loss function and after applying the cuts post-training, as a function of the target efficiency for subnets in ensemble of one-to-one linear networks trained using the SUSY dataset.
7 Discussion
Event selections using pass/fail cuts on input features are ubiquitous in high-energy physics, from phase space restrictions in Monte Carlo Event generators to hardware triggers and offline analysis. Despite their prevalence, cut optimization is often left to ad hoc procedures, as opposed to more optimization of more complicated discriminants based on neural network or BDT outouts.
To fill this gap in optimization tools, we present the CABIN approach to learning optimized cuts for event classification, where the cut values are represented by biases in a type of linear neural network. The CABIN technique uses gradient-based learning and is aware of useful HEP metrics such as signal and background efficiency. The method can self-regulate cuts across working points to deliver both good performance and robustness against statistical noise. An implementation of this approach is available in the cabin python package [cabin].
Extensions of this approach can be used to optimize cuts based on significance metrics [ATL-PHYS-PUB-2020-025], possibly with additional requirements to ensure signal yield large enough to enable discovery or limit setting. Another foreseen extension will be ensuring the smooth evolution cuts as functions of particle or event properties that cause correlated changes in input features. For example, when training a classifier for particle identification, the input features may vary with particle energy, and will likely evolve monotonically, though possibly in different ways for signal and background. Contractive terms in the loss function can ensure that cuts on input features for neighboring energy bins have similar values, consistent with the expected smooth evolution of the features themselves.
Finally, as suggested in Ref. [Watts:2024wes], a CABIN network can be embedded as one step in a chain of differentiable analysis selections. This provides a path towards full-chain optimization of cut-based event generation, triggering, and analysis pre-selection, followed by final optimization using fully connected networks, all within the same minimization task.
Acknowledgments
We would like to thank Earl Almazan, Scott Phillips, and Giordon Stark for helpful discussions on the cabin approach. This work was supported by the U.S. Department of Energy (DOE) Office of High Energy Physics under Grant No. DE-SC0010107.
[…]